OpenStudy (anonymous):

Write a polynomial function of least degree that has given zeros. 2i, -2i

OpenStudy (jordanloveangel):

Reverse solve for x for the zeroes. x = -2 x + 2 = 0 x = -1 x + 1 = 0 x = 1 x - 1 = 0 f(x) = (x + 2)(x + 1)(x - 1) Can you multiply it out?

OpenStudy (jordanloveangel):

f(x) = (x+2)(x+1)(x-1) Multiply it out See that putting -1,-1 or 1 in the abuv function gives 0

OpenStudy (jordanloveangel):

hop these help

OpenStudy (anonymous):

I don't get it

OpenStudy (jdoe0001):

\(\large { 2i, -2i\quad \begin{cases} x=2i\to x-2i=0\to &(x-2i)=0\\ x=-2i\to x+2i=0\to &(x+2i)=0 \end{cases} \\ \quad \\ (x-2i)(x+2i)=\textit{original polynomial} }\)